OpenStudy (anonymous):

3/2x - 15 = -12 solving equations containing fractions

OpenStudy (anonymous):

Is it \[\frac{ 3x }{ 2 } - 15 = -12\]

OpenStudy (anonymous):

If it is, then here's the procedure: \[\frac{ 3x }{2 } = -12 + 15\] \[3x = 3 \times 2\] You should be able to do it now, aye?

OpenStudy (anonymous):



OpenStudy (anonymous):

If it's \[\frac{ 3 }{ 2x } - 15 = -12\] Then, here is the procedure: \[\frac{ 3 }{ 2x } = -12 + 15\] \[3 = 2x \times 3\] \[2x = 1\] And therefore, \[x = \frac{ 1 }{2 }\]